OpenStudy (anonymous):

What is curium-240 Decays By alpha Emission?

thomaster (thomaster):

\[\begin{matrix}240 \\ 96\end{matrix}Cm\rightarrow\begin{matrix}236 \\ 94\end{matrix}Pu+\begin{matrix}4 \\ 2\end{matrix}He\]